OpenStudy (anonymous):

If f(x) = x^2 and g(x) = 1 over 2x + 3 , find g(f(-1)). Help!!! pls

OpenStudy (anonymous):

help pls

OpenStudy (anonymous):

what is \(f(-1)\)?

OpenStudy (anonymous):

is it -1 ?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

I think it should be 1/2(-1)^2 +3

OpenStudy (anonymous):

\[(-1)^2 \]

OpenStudy (anonymous):

like it asks If f(x) = x^2 and g(x) = 1 over 2x + 3 , find g(f(-1)). A.-1 B. 1/25 C.1/5 D. 1

OpenStudy (anonymous):

First find \(f(-1)\).

OpenStudy (anonymous):

isn't f (-1) -1?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

Okay, I guess I have to do it then....

OpenStudy (anonymous):

\[ f(-1) = (-1)^2=(-1)(-1)=1 \]

OpenStudy (anonymous):

Can you comprehend this?

OpenStudy (anonymous):

Do I need to start drawing pictures?

OpenStudy (anonymous):

no i get it y is 1 now

OpenStudy (anonymous):

So do you see that \(g(f(-1)) = g(1)\)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Can you solve \(g(1)\)?

OpenStudy (anonymous):

don't you plug it in the 1 over 2x+3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait if u plug that in wouldn't it be 2 over 2x+3???

OpenStudy (anonymous):

??

OpenStudy (anonymous):

u there

OpenStudy (anonymous):

You change \(x\) to be \((1)\)

OpenStudy (anonymous):

then simplify

OpenStudy (anonymous):

so would the answer be 1 over 5 ?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

can u pls just tell me the answer its 3 am n i start school at 6 pls

OpenStudy (anonymous):

??????????????????

OpenStudy (anonymous):

it is \(1/5\)